OpenStudy (anonymous):

\[\int_{0}^{1} \sqrt(1+\frac{x^2}{4-x^2})\]

OpenStudy (anonymous):

Is it?? \[\int\limits_{0}^{1} \sqrt{1+\frac{x^2}{2-x^2}}dx\]

OpenStudy (anonymous):

sorry no it's 4-x^2

OpenStudy (anonymous):

So is it? \[\int\limits_{0}^{1} \sqrt{1+\frac{x^2}{4-x^2}}\]

OpenStudy (anonymous):

Take the LCM of the denominator... \[\frac{4 - x^2 + x^2}{4 - x^2} = \frac{4}{4 - x^2}\]

OpenStudy (campbell_st):

doesn't this become \[\int\limits_{0}^{1} \sqrt{\frac{4}{4 - x^2}} dx\]

OpenStudy (campbell_st):

which can be written as \[2\int\limits_{0}^{1}\sqrt{\frac{1}{4 -x^2}} dx\]

OpenStudy (campbell_st):

and \[2\int\limits_{0}^{1}\frac{1}{\sqrt{4 -x^2}} dx\]

OpenStudy (anonymous):

Use the following formula: \[\int\limits_{}^{}\frac{1}{\sqrt{a^2 - x^2}}.dx = \sin^{-1}\frac{x}{a} + C\] Here a = 2..

OpenStudy (anonymous):

got it, thanks!

OpenStudy (anonymous):

Welcome dear..